Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Message ID: 241     Entry time: Fri Mar 8 09:20:29 2019
Author: Drive Team- Ino&Thierry&Armand 
Type: Telescope Operations 
Category: General 
Subject: High speed drive tests 
Status:  
  • On Thrusday :
    • new movements at 45% from -90 to 270 have been done.
    • No possible movement at 60% because the FW and UPS issues are not solved yet
      • Ino will back on site when HITACHI team will be on site for FW maintenance
        • Tests at 60% will be done to validate (or not) the correction applied by Hitachi team.
ELOG V3.1.5-fc6679b